# Patherea: Cell Detection and Classification for the 2020s

### v1

@@ 6 Experiments
In this section we report results on two established public datasets for point-based cell detection and classification - Lizard (Graham et al., 2021) and BRCA-M2C (Abousamra et al., 2021) in Section 6.2. Additionally, we also report results on BCData (Huang et al., 2020) in A.3. We report results on our newly introduced Patherea dataset in Section 6.3. We compare the proposed Patherea-P2P method introduced in Section 3 against the recent DETR-based approaches ACFormer (Huang et al., 2023b) and PGT (Huang et al., 2023a), as well as against more traditional approaches with an intermediate representation(s). For the traditional approaches, we compared against MCSpatNet (Abousamra et al., 2021), which additionally utilizes spatial context information and against our implementation of the approach from (Xie et al., 2018) - named Patherea-FCRN, which was modernized with ResNet-34 backbone and added support for cell classification, mostly following (Lee et al., 2021). We used the official code for ACFormer, PGT and MCSpatNet, including their publicly available pre-trained weights, when available for Lizard (A.1) or BRCA-M2C (A.2) datasets. We will publicly release the modernized re-implementation of FCRN (Xie et al., 2018) as no official code was released with the paper. F1 scores for Lizard, BRCA-M2C and Patherea datasets have been reported using the evaluation protocol described in Section 5.
@@ 6.1 Implementation Details
Patherea-P2P and FCRN (Xie et al., 2018) methods were implemented in PyTorch, while the official code and parameters were used to reproduce MCSpatNet (Abousamra et al., 2021), ACFormer (Huang et al., 2023b) and PGT methods (Huang et al., 2023a) on selected datasets. Patherea-P2P proposal candidates were initialized with and a Feature Pyramid Network (FPN) (Lin et al., 2017) was used for ConvNext (Liu et al., 2022) and ViT (Dosovitskiy et al., 2021) backbones with number of features set to 256 for ConvNext-B and 768 for ViT-B. Only higher-resolution features at level 2 of the FPN were used as an input to and . Lightweight regression and classification heads were attached with two ResNet blocks (He et al., 2016) and two convolutions for one-to-one and one-to-many hybrid Hungarian matching.
We use as a weight term for the pixel distance in the Hungarian matching. In a hybrid Hungarian matching setup, was used for public datasets reported in Section 6.2, while was used for the Patherea dataset, due to the increased resolution of the dataset. The influence of parameter is investigated in Section 6.5. A class weighting term in was set to 0.5 for the background class and 10 for all the foreground classes . Regression loss weight in and was set to , while in a combined loss was set to 0.5. Threshold on confidence in inference was set to 0.9 for all classes for Lizard(Graham et al., 2021) and BRCA-M2C (Abousamra et al., 2021) datasets, while 0.5 was used for Patherea dataset.
ImageNet pre-trained weights were used to initialize the backbone, except when pathology foundational model (Nechaev et al., 2024) was specifically mentioned. Both, Patherea-P2P and FCRN methods were trained for 1000 epochs on public datasets (Section 6.2), while 100 epochs were used for the Patherea dataset in Section 6.3. This is due to the much larger Patherea dataset, which makes training competing methods inefficient for larger training cycles. The batch size was set to 16 and AdamW (Loshchilov and Hutter, 2019) was used as an optimizer with a learning rate set to , weight decay to and cosine annealing scheduler, with a linear warm-up in all experiments.
@@ 6.2 Public Datasets
Lizard: Lizard dataset was introduced in (Graham et al., 2021) and consists out of 6 different public datasets of colon cancer, where 291 images were extracted at x20 objective magnification. The Lizard dataset was primarily developed for nuclear instance segmentation and classification, but also provided cell center locations which can be utilized to develop and evaluate point-based approaches. The dataset was mostly automatically labeled, with a HoVer-Net (Graham et al., 2019) trained on existing public data and used as an initial segmentation result. The segmentation results were later refined by a pathologists and a model was re-trained. Similarly, cell class refinement was augmented with pathologists-in-the-loop. This semi-automatic approach enabled the annotation of 495,179 cells of 6 different classes (epithelial, lymphocyte, plasma, neutrophil, eosinophil, connective).
@@
We report the results using the 5x re-training strategy to demonstrate real-world robustness and reproducibility. We report the average of the F1 scores across the 5 training runs. The official training code and recipes were used to re-train the models, except for the MCSpatNet and our re-implementation of the FCRN method, which did not report the results on the Lizard dataset. The checkpoint was selected using the best results on the validation set. We also report results using the officially provided pre-trained weights on Lizard for ACFormer and PGT in A.1, where the improvements of our model are even more significant.
The best results are achieved using our proposed Patherea-P2P approach, using the ConvNext backbone. We hypothesize that the dataset of only 291 is too small for the ViT backbone to generalize and learn the inductive biases, despite ImageNet initialization. Another reason could be low resolution of input images, scanned at 20x objective magnification and compressed, which basically removes most of the cell morphology information. We slightly improved ViT results by using a pathology foundational model† (Section 6.5). We also noticed that density-based approaches MCSpatNet and FCRN performed significantly worse, mostly due to inability to detect eosinophil and neutrophil classes. These two classes are in low abundance (e.g., 3,604 and 4,824 labels in comparison with more than 100k for epithelial, lymphocyte and connective).
BRCA-M2C: This breast cancer dataset was introduced with the MCSpatNet (Abousamra et al., 2021) and consists of 120 patches belonging to 113 patients, collected from TCGA. The dataset was collected at 20x objective magnification with inflammatory (lymphocyte), epithelial and stromal cells being labeled as point-based annotations. The ground truth radious was used for evaluation. We report results in Table 4.
@@
We report the results using the 5x re-training strategy and report the average F1 scores across the 5 training runs. Our proposed approach again achieved the best results using the ConvNext backbone, due to the small size of the BRCA-M2C dataset. The performance of the ViT-based backbone improved when the pathology foundational model† was used for pre-training. We also report results using the official weights for ACFormer and PGT in A.2.
@@ 6.3 Patherea Dataset
We evaluate the proposed Patherea-P2P method on our newly proposed Patherea dataset introduced in Section 4. We split the Patherea dataset into 3 folds across different patients in a random manner and perform 3-fold cross-validation for all the approaches. The splits across the folds are released alongside the Patherea dataset. The normal cells were excluded from the analysis, as only 80 cells were labeled across all parts of the Patherea dataset. We used a fixed training schedule of 100 epochs for all the methods and report the average across the folds at 100 epochs. We report the results across different parts of the Patherea dataset - LNET in Table 5, GNET in Table 6 and Breast in Table 7.
Overall, Patherea-P2P approach outperforms all of the competing approaches by a significant margin (e.g., 4-13%) in comparison with the second best approach across different Pathera datasets. The improvement is more significant with less abundant classes of others positive and others negative classes. Interestingly, density-estimation-based approaches MCSpatNet and FCRN outperformed recently proposed DETR-based approaches.
| MCSpatNet | 0.790 | 0.734 | 0.190 | 0.446 | 0.540 | 1d18h/1 |
| ACFormer | 0.826 | 0.781 | - | - | 0.402 | 1d18h/4 |
| PGT | 0.789 | 0.758 | - | - | 0.387 | 1d19h/4 |
| FCRN | 0.787 | 0.750 | 0.245 | 0.467 | 0.562 | 4h/1 |
| ours (ViT) | 0.832 | 0.790 | 0.273 | 0.572 | 0.617 | 2.1h/1 |
| ours (CNN) | 0.828 | 0.789 | 0.318 | 0.559 | 0.624 | 1.5h/1 |
@@
MCSpatNet 0.771/0.690 0.680/0.696 0.315/- 0.606/0.511 0.593/0.474 ACFormer 0.761/0.717 0.657/0.668 -/- - /- 0.378 PGT 0.794/0.770 0.713/0.742 -/- -/- 0.377/0.378 FCRN 0.785/0.768 0.736/0.762 0.326/0.025 0.586/0.533 0.608/0.522 ours (ViT) 0.804/0.776 0.748/0.763 0.341/0.143 0.640/0.585 0.633/0.567 ours (CNN) 0.752/0.714 0.698/0.722 0.319/0.169 0.607/0.531 0.594/0.534
We also report an approximate training time per-fold on LNET dataset in Table 5. Training time of our proposed Patherea-P2P method is multi-fold faster, regardless of the chosen backbone. Note that we didn’t include pre-processing time needed to create intermediate representations for MCSpatNet and FCRN. We also only report the prompt fine-tuning training time for PGT.
@@ 6.4 Qualitative Results
Qualitative results for the Patherea-P2P model are presented in Figure 4. First, we notice that proposal candidates are successfully filtered down with the Hungarian one-to-one matching and confidence thresholding (0.5), such that multiple detection occur rarely. This also holds true in most of the extreme cases, where cell size is significantly larger (e.g., LNET-4, GNET-3). Cells are successfully detected in dense regions (e.g., LNET-2, GNET-2), as well as when cell morphology is less apparent (e.g., GNET-5, Breast-4). There are some multiple detections present when cells are significantly overlapped and clustered (e.g., Breast-5). One could utilize the non-maxima suppression to further fine-tune the predictions, which would result in the introduction of the post-processing step and break the fully end-to-end learning-based paradigm.
Labeling cells directly from WSI, replicating the real-world diagnostics represents a significant challenge in comparison with directly labeling small patches. Certainly, some cells will be missed or different cut-off thresholds will be used by different pathologists to label the specific cell. We notice some detections, that were not labeled by expert pathologists (e.g., LNET-4, GNET-3). We can also notice (Breast-4) how the model fits the pathologist that was labeling the data. Breast samples represent the same samples being labeled twice by two different pathologists. We notice, that they mostly disagree in border classes of others positive and others negative.
In Figures 5 and 6 we present qualitative results for Lizard (Graham et al., 2021) and BRCA-M2C (Abousamra et al., 2021) datasets. We notice that resolution is significantly lower with significantly less cell morphology visible to differentiate different cell types. We notice a confident detection of well-representative classes and similarly some spurious detections in regions where no labels are present.
@@ 6.5 Ablation Studies
Hybrid Hungarian Matching: In Table 8 we evaluate the influence of the parameter which is used to select the number of candidates each ground truth target is matched against for an additional one-to-many loss in Patherea-P2P, as presented in Section 3.2. We demonstrate the effectiveness of the proposed hybrid matching, with relative performance gains of up to 7% for sufficiently represented classes and up to almost 25% for less abundant classes.
@@
We notice that performance on Patherea-LNET dataset is improving with increasing and starts to saturate between and . We used as default for Patherea dataset in Section 6.3, while was used for Lizard and BRCA-M2C datasets in Section 6.2 due to only x20 objective magnification used.
@@
Foundational Model: In Table 9 we report the average F1 performance on Lizard (Graham et al., 2021), BRCA-M2C (Abousamra et al., 2021) and Patherea-LNET datasets when using different model training strategies. For Lizard and BRCA-M2C we report the across different classes with a 5x re-training strategy, directly comparable with results reported in Tables 3 and 4. Similarly to Table 5, we report across the 3 folds for Patherea-LNET dataset.
@@
We evaluated three different training strategies for training Patherea-P2P using the ViT backbone. i) With the Scratch strategy, we trained the Patherea-P2P model from scratch (backbone, ViT-Adapter, heads). ii) The 2nd approach, fine-tuning from ImageNet initialized ViT backbone represents the default approach used to report results in Sections 6.2 and 6.3. iii) Lastly, we utilized the recently released Hibou family of open-source foundational models for pathology (Nechaev et al., 2024). For ImageNet, we fine-tuned the whole Patherea-P2P architecture, while for Hibou, we only fine-tined ViT-Adapter (Chen et al., 2023) and heads and kept the ViT backbone frozen.
We notice that training from scratch significantly reduces the performance on Lizard and BRCA-M2C datasets, while the drop on the Patherea-LNET dataset is smaller. We hypothesize that this is due to significantly smaller size of Lizard and BRCA-M2C datasets in terms of the actual number of training samples. We also notice a 3-5% improvement over ImageNet when using a foundational model on Lizard and BRCA-M2C datasets. We were able to improve upon ConvNext-based Patherea-P2P results on BRCA-M2C, reported in Table 4, when using a foundational model. No improvement was observed on Patherea-LNET dataset when using a pre-trained Hibou foundational model. We hypothesize that this is due to the training data used in Hibou, where mostly H&E samples were used for training, same staining as used in Lizard and BRCA-M2C datasets. We expect to observe similar gains when foundational models start to appear, that were predominately trained on Ki-67 staining.
@@ 7 Conclusion
Cell detection and classification represents an essential tool to perform diagnostics in various pathology workflows. Such diagnostics workflows are extremely time-consuming for pathologists, which in practice results in performing them less accurately, leading to lower interobserver concordance and reproducibility. Automatisation of such workflows requires a significant amount of labeled data, which for the task of cell detection and classification is most efficiently obtained as point-based annotations. There is also a lack of approaches that can effectively utilize such point-based annotations and can be easily deployed in clinical practice. Ideally, the models can effectively utilize pre-trained foundational models to generalize across different cell detection and classification tasks with limited training data and/or enable efficient re-training or fine-tuning on newly collected data.
In this work, we introduced a Patherea framework that directly addresses the development of AI models in a clinical setting. First, we collected the largest fully manually labeled point-based cell detection and classification dataset for a particular clinical task of Ki-67 proliferation index estimation in Neuroendocrine tumors (NETs) and breast cancer. The annotation process was designed to mimic the ideal clinical setting, where pathologists would be manually counting the cells for the estimation of the Ki-67 index, instead of the less accurate "eye-balling" technique.
Secondly, we introduced a new approach for an automated cell detection and classification that can directly utilize point-based annotations, without the need for intermediate representations. We proposed a more efficient approach for the target-candidate association that can effectively utilize multiple (good) proposal candidates for a particular cell. We reported state-of-the-art results on existing large-scale public datasets and demonstrated that the newly proposed Patherea dataset presents a significant challenge for existing approaches, while the proposed Patherea-P2P achieving significantly better results. We also demonstrated that the proposed architecture can be effectively used to utilize existing large-scale unlabeled data to pre-train large-scale foundational models and apply them to specific tasks with a limited amount of labeled data.
We hope that the release of the Patherea framework will facilitate the development of new approaches that can effectively handle imbalanced datasets and distinguish challenging cell type classes introduced with our dataset. Comparing different approaches with reliable metrics is important, and we have revised the established scoring protocols for point-based cell detection and classification and proposed a more accurate approach.
Finally, in this work, we only addressed the specific problem of point-based cell detection and classification, without addressing the actual clinical problem of Ki-67 proliferation index estimation. In our future work, we plan to apply the proposed approaches to a large cohort of Ki-67 stained samples that have slide-level Ki-67 index estimations provided by multiple pathologists and compare that against the automatically estimated Ki-67 index that was calculated based on the approach developed in this work.

### v2

@@ 6 Experiments
In this section we report results on two established public datasets for point-based cell detection and classification - Lizard (Graham et al., 2021) and BRCA-M2C (Abousamra et al., 2021) in Section 6.2. We report results on our newly introduced Patherea dataset in Section 6.3. We compare the proposed Patherea-P2P method introduced in Section 3 against the recent DETR-based approaches ACFormer (Huang et al., 2023b) and PGT (Huang et al., 2023a), as well as against more traditional approaches with an intermediate representation(s). For the traditional approaches, we compared against MCSpatNet (Abousamra et al., 2021), which additionally utilizes spatial context information. We also compared against our implementation of the approach from (Xie et al., 2018), named FCRN++. This implementation was modernized with a ResNet-34 backbone and added support for cell classification, mostly following (Lee et al., 2021). We used the official code for ACFormer, PGT and MCSpatNet. When available, we also utilized their publicly available pre-trained weights for the Lizard (B.1) or BRCA-M2C (B.2) datasets. Additionally, we also report results on BCData (Huang et al., 2020) in B.3. F1 scores for Lizard, BRCA-M2C and Patherea datasets have been reported using the evaluation protocol described in Section 5.
@@ 6.1 Implementation Details
Patherea-P2P and FCRN++ methods were implemented in PyTorch, while the official code and parameters were used to reproduce MCSpatNet (Abousamra et al., 2021), ACFormer (Huang et al., 2023b) and PGT methods (Huang et al., 2023a) on selected datasets. Patherea-P2P proposal candidates were initialized with and a Feature Pyramid Network (FPN) (Lin et al., 2017) was used for ConvNext (Liu et al., 2022) and ViT (Dosovitskiy et al., 2021) backbones with number of features set to 256 for ConvNext-B and 768 for ViT-B. Only higher-resolution features at level 2 of the FPN were used as an input to and . Lightweight regression and classification heads were attached with two ResNet blocks (He et al., 2016) and two convolutions for one-to-one and one-to-many hybrid Hungarian matching. The models were trained and evaluated in an HPC environment using Nvidia A100 40GB GPUs. We provide training and inference runtime analysis for different approaches in B.7.
We use as a weight term for the pixel distance in the Hungarian matching. In a hybrid Hungarian matching setup, was used for public datasets reported in Section 6.2, while was used for the Patherea dataset, due to the increased resolution of the dataset. The influence of parameter is investigated in Section 6.5. A class weighting term in was set to 0.5 for the background class and 10 for all the foreground classes . Regression loss weight in and was set to , while in a combined loss was set to 0.5. Threshold on confidence in inference was set to 0.9 for all classes for Lizard (Graham et al., 2021) and BRCA-M2C (Abousamra et al., 2021) datasets, while 0.5 was used for the Patherea dataset.
All models were initialized with ImageNet pre-trained weights for the backbone network, except where the pathology foundation model (Nechaev et al., 2024; Chen et al., 2024) was explicitly employed. Training was conducted for 1000 epochs on public datasets (Section 6.2) for both Patherea-P2P and FCRN++ methods. For experiments on the Patherea dataset (Section 6.3), 100 epochs were used given the substantially larger dataset size, which would render extended training cycles computationally intractable for competing methods. All experiments employed a batch size of 16 and utilized the AdamW optimizer (Loshchilov and Hutter, 2019) with a learning rate of , weight decay of , and cosine annealing scheduling with a linear warm-up.
@@ 6.2 Public Datasets
Lizard: Lizard dataset was introduced in (Graham et al., 2021) and consists out of 6 different public datasets of colon cancer, where 291 images were extracted at x20 objective magnification. The Lizard dataset was primarily developed for nuclear instance segmentation and classification, but also provided cell center locations which can be utilized to develop and evaluate point-based approaches. The dataset was mostly automatically labeled, with a HoVer-Net (Graham et al., 2019) trained on existing public data and used as an initial segmentation result. The segmentation results were later refined by a pathologist and a model was re-trained. Similarly, cell class refinement was augmented with pathologists-in-the-loop. This semi-automatic approach enabled the annotation of 495,179 cells of 6 different classes (epithelial, lymphocyte, plasma, neutrophil, eosinophil, connective).
@@
We report the results using the 5x re-training strategy to demonstrate real-world robustness and reproducibility. We report the average of the F1 scores across the 5 training runs. The official training code and recipes were used to re-train the models, except for the MCSpatNet and our re-implementation of the FCRN++ method, which did not report the results on the Lizard dataset. The checkpoint was selected using the best results on the validation set. We also report results using the officially provided pre-trained weights on Lizard for ACFormer and PGT in B.1, where the improvements of our model are even more significant.
The best results are achieved using our proposed Patherea-P2P approach, using the ConvNext backbone. We hypothesize that the dataset of only 291 is too small for the ViT backbone to generalize and learn the inductive biases, despite ImageNet initialization. Another reason could be the low resolution of input images, scanned at 20x objective magnification and compressed, which basically removes most of the cell morphology information. We slightly improved ViT results by using a pathology foundation model† (Section 6.5). We also noticed that density-based approaches MCSpatNet and FCRN++ performed significantly worse, mostly due to inability to detect eosinophil and neutrophil classes. These two classes are in low abundance (e.g., 3,604 and 4,824 labels in comparison with more than 100k for epithelial, lymphocyte and connective).
BRCA-M2C: This breast cancer dataset was introduced with the MCSpatNet (Abousamra et al., 2021) and consists of 120 patches belonging to 113 patients, collected from TCGA. The dataset was collected at 20x objective magnification with inflammatory (lymphocyte), epithelial and stromal cells being labeled as point-based annotations. The ground truth radius was used for evaluation. We report results in Table 4.
@@
We report the results using the 5x re-training strategy and report the average F1 scores across the 5 training runs. Our proposed approach again achieved the best results using the ConvNext backbone, due to the small size of the BRCA-M2C dataset. The performance of the ViT-based backbone improved when the pathology foundation model† was used for pre-training. We also report results using the official weights for ACFormer and PGT in B.2.
@@ 6.3 Patherea Dataset
We evaluate the proposed Patherea-P2P method on our newly proposed Patherea dataset introduced in Section 4. We split the Patherea dataset into 3 folds across different patients in a random manner and perform 3-fold cross-validation for all the approaches. The splits across the folds are released alongside the Patherea dataset. We used a fixed training schedule of 100 epochs for all the methods and report the average across the folds at 100 epochs. We report the results across different parts of the Patherea dataset - LNET in Table 5, GNET in Table 6, and Breast in Table 7.
Overall, the Patherea-P2P approach outperforms all of the competing approaches by a significant margin (e.g., 4-13%) in comparison with the second best approach across different Patherea datasets. The improvement is more significant with less abundant classes of others positive and others negative classes. Interestingly, density-estimation-based approaches MCSpatNet and FCRN++ outperformed recently proposed DETR-based approaches.
| MCSpatNet | 0.790 | 0.734 | 0.190 | 0.446 | 0.540 |
| ACFormer | 0.826 | 0.781 | - | - | 0.402 |
| PGT | 0.789 | 0.758 | - | - | 0.387 |
| FCRN++ | 0.787 | 0.750 | 0.245 | 0.467 | 0.562 |
| ours (ViT) | 0.832 | 0.790 | 0.273 | 0.572 | 0.617 |
| ours (CNN) | 0.828 | 0.789 | 0.318 | 0.559 | 0.624 |
@@
MCSpatNet 0.771/0.690 0.680/0.696 0.315/- 0.606/0.511 0.593/0.474 ACFormer 0.761/0.717 0.657/0.668 -/- - /- 0.378 PGT 0.794/0.770 0.713/0.742 -/- -/- 0.377/0.378 FCRN++ 0.785/0.768 0.736/0.762 0.326/0.025 0.586/0.533 0.608/0.522 ours (ViT) 0.804/0.776 0.748/0.763 0.341/0.143 0.640/0.585 0.633/0.567 ours (CNN) 0.752/0.714 0.698/0.722 0.319/0.169 0.607/0.531 0.594/0.534
@@ 6.4 Qualitative Results
Qualitative results for the Patherea-P2P model are presented in Figure 3. First, we notice that proposal candidates are successfully filtered down with the Hungarian one-to-one matching and confidence thresholding (0.5), such that multiple detections occur rarely. This also holds true in most of the extreme cases, where cell size is significantly larger (e.g., LNET-4, GNET-3). Cells are successfully detected in dense regions (e.g., LNET-2, GNET-2), as well as when cell morphology is less apparent (e.g., GNET-5, Breast-4). There are some multiple detections present when cells are significantly overlapped and clustered (e.g., Breast-5). One could utilize the non-maxima suppression to fine-tune further the predictions, which would result in the introduction of the post-processing step and break the fully end-to-end learning-based paradigm. Qualitative comparative analysis against other approaches is presented in Figure 5.
Labeling cells directly on whole slide images (WSIs), as done in real-world diagnostic workflows, poses a substantially greater challenge than annotating isolated image patches. Variability arises not only from missed annotations but also from subjective differences in the thresholds applied by individual pathologists when labeling specific cell types. For instance, we observe model-predicted detections that were not annotated by experts (e.g., LNET-4, GNET-3), suggesting possible under-labeling or inter-observer variability. In the Breast-4 sample, the model appears to align closely with the labeling behavior of the specific pathologist, highlighting the model’s ability to learn annotator-specific patterns. Notably, breast samples were independently labeled twice by different pathologists, and their annotations often diverge in borderline categories — particularly between others positive and others negative — emphasizing the subjectivity involved in such classifications.
In Figures 6 and 7 we present qualitative results for Lizard (Graham et al., 2021) and BRCA-M2C (Abousamra et al., 2021) datasets. We notice that the resolution is significantly lower with significantly less cell morphology visible to differentiate different cell types. We notice a confident detection of well-representative classes and similarly some spurious detections in regions where no labels are present.
@@ 6.5 Ablation Studies
Hybrid Hungarian Matching: In Table 8 we evaluate the influence of the parameter which is used to select the number of candidates each ground truth target is matched against for an additional one-to-many loss in Patherea-P2P, as presented in Section 3.2. We demonstrate the effectiveness of the proposed hybrid matching, with relative performance gains of up to 7% for sufficiently represented classes and up to 24% for less abundant classes.
@@
We observe that the performance on the Patherea-LNET dataset improves with increasing values of , with performance gains beginning to plateau between and . Accordingly, we set as the default for the Patherea dataset in Section 6.3. In contrast, for the Lizard and BRCA-M2C datasets, a lower value of was used in Section 6.2, reflecting the lower resolution associated with the 20 objective magnification used in those datasets.
@@
Foundation Model: In Table 9 we report the average F1 performance on Lizard (Graham et al., 2021), BRCA-M2C (Abousamra et al., 2021) and Patherea-LNET datasets when using different model training strategies. For Lizard and BRCA-M2C we report the across different classes with a 5x re-training strategy, directly comparable with results reported in Tables 3 and 4. Similarly to Table 5, we report across the 3 folds for the Patherea-LNET dataset.
@@
We evaluated three different training strategies for training Patherea-P2P using the ViT backbone. With the Scratch strategy, we trained the Patherea-P2P model from scratch (backbone, ViT-Adapter, heads). The 2nd approach, fine-tuning from ImageNet initialized ViT backbone represents the default approach used to report results in Sections 6.2 and 6.3. Lastly, we utilized the recently released Hibou family of open-source foundation models for pathology (Nechaev et al., 2024). For ImageNet, we fine-tuned the whole Patherea-P2P architecture, while for Hibou, we only fine-tuned ViT-Adapter (Chen et al., 2023) and heads and kept the ViT backbone frozen.
We notice that training from scratch significantly reduces the performance on the Lizard and BRCA-M2C datasets, while the drop on the Patherea-LNET dataset is smaller. We hypothesize that this is due to the significantly smaller size of the Lizard and BRCA-M2C datasets in terms of the actual number of training samples. We also notice a 3-5% improvement over ImageNet when using a foundation model on Lizard and BRCA-M2C datasets. We were able to improve upon ConvNext-based Patherea-P2P results on BRCA-M2C, reported in Table 4, when using a foundation model. No improvement was observed on the Patherea-LNET dataset when using a pre-trained Hibou foundation model.
We hypothesize that this is due to the training data used in Hibou, where around 80% of the training samples are H&E stains, with the other 20% represented by other stains, which are not specified. This is common with the current publicly available foundation models, as immunohistochemistry samples are not widely available. We additionally confirm the improved results on Lizard and BRCA-M2C using the UNI foundation model (Chen et al., 2024) in Table 14. We expect to observe similar gains on the Patherea dataset with foundation models, specifically trained on Ki-67 samples.
@@ 7 Conclusion
In this work, we presented Patherea, a comprehensive framework for developing clinically applicable AI models for point-based cell detection and classification. Diagnostic workflows involving immunohistochemistry, such as Ki-67 proliferation index estimation, remain labor-intensive and subject to interobserver variability. Automating these tasks is of critical importance, yet it is hindered by the scarcity of large, high-quality annotated datasets and the limited capacity of existing methods to effectively utilize point-based annotations—the most efficient annotation modality in clinical practice.
To address this gap, we curated the largest fully manually annotated dataset with point-based labels for Ki-67 estimation in Neuroendocrine tumors (NETs) and breast cancer. Importantly, the annotation protocol was designed to closely mirror real-world clinical workflows: pathologists performed cell annotations while having full access to the entire whole slide image (WSI), rather than being constrained to isolated image patches. This approach ensures that annotations reflect the contextual decision-making process used in routine diagnostics. The resulting dataset thus provides both a realistic benchmark for evaluating automated methods and a high-quality resource for training models that are aligned with actual clinical practice.
We introduced a novel approach for cell detection and classification that learns directly from point annotations without requiring intermediate representations such as bounding boxes or segmentation masks. Our proposed method includes an improved target-candidate association strategy that considers multiple high-quality candidates for each annotated point, thereby enhancing robustness and precision. The resulting model, Patherea-P2P, achieves state-of-the-art performance on existing large-scale public benchmarks and demonstrates superior generalization on the newly introduced Patherea dataset, which poses substantial challenges due to its clinical complexity.
Additionally, we demonstrated that the proposed architecture is well-suited for leveraging large-scale unlabeled data through foundation model pre-training. This enables effective transfer learning and fine-tuning with limited labeled data, which is critical for real-world clinical applications. We also revisited the evaluation protocols commonly used in point-based detection and classification tasks, proposing a revised and more accurate scoring scheme to ensure reliable and meaningful comparisons across methods.
While the focus of this study is the technical development of point-based cell detection and classification methods, our ultimate goal is to support accurate and reproducible estimation of clinically relevant biomarkers, such as the Ki-67 index. As part of future work, we plan to apply the proposed framework to large cohorts of Ki-67 stained samples annotated with slide-level index scores from multiple pathologists. This will enable rigorous evaluation of the clinical validity of our approach and further establish its utility in diagnostic pathology.